OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10?

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

solve for y, and look at the coefficient of x.

OpenStudy (anonymous):

can you help me agdgdgdgwngo?

jimthompson5910 (jim_thompson5910):

\[\Large 5x + 2y = 10\] \[\Large 2y = 10-5x\] \[\Large 2y = -5x+10\] \[\Large y = \frac{-5x+10}{2}\] \[\Large y = \frac{-5x}{2}+\frac{10}{2}\] \[\Large y = -\frac{5}{2}x+5\] So the equation is now in slope intercept form y = mx+b where m is the slope and (0, b) is the y-intercept. In this case, we see that \[\Large m = -\frac{5}{2}\] So the slope of the line is \[\Large -\frac{5}{2}\]